A rod of resistive material of length L is connected to a battery with potential difference 12 V. The resistance per unit length of the rod varies as ρ(x)=12x ohm/m where x is the distance from one end of the rod.If the current in the resistive rod transfer energy to thermal energy at the rate of 600W then the length of rod will be
Ans 20cm
1
Expert's answer
2020-02-03T14:07:39-0500
"R=\frac{U^2}{P}=\frac{12^2}{600}=0.24\Omega"

"R=12\frac{L^2}{2}=6L^2=0.24"

"L=0.2\ m=20\ cm"